OpenStudy (anonymous):

(2(x-3))/(6sqrt ofx(x-3))^2

OpenStudy (anonymous):

i need this formula squared please

OpenStudy (anonymous):

*equation

OpenStudy (anonymous):

2x-3/6 sqrt of x (x-3)

OpenStudy (anonymous):

You have to to square top and bottom separately.

OpenStudy (anonymous):

\[(2(x-3)/(6\sqrt{x(x-3)})^{2}=1/18\]

OpenStudy (anonymous):

o ok i thought i have to do it as a whole ok thamks

OpenStudy (anonymous):